A 8-quart radiator contains a 60% antifreeze solution. How much of the solution needs to be drained out and replaced with pure antifreeze in order to raise the solution to 80% antifreeze?
Answer by mananth(16949)

you have 60% solution
you have to add 100% antifreeze to get 80% solution
so we find out how much antifreeze is required to raise ti 80%
This quantity is to be replaced
percent ---------------- quantity
Antifreeze 100 ----------------x quarts
solution 60 ----------------5-x quarts
Mixture 80.00% ---------------- 5

100x+60(5-x)=80*5

100x+300-60x=400
100x-60 x=400-300
40x= 100
/ 40
x=2.5 quarts 100.00% Antifreeze
2.5 quarts 60.00% solution
2.5 quarts has to be removed
m.ananth@hotmail.ca